A pickup truck has a width of 61.7 in. If it is traveling north at 32 m/s through a magnetic field wth vertical component of 36 μT, what magnitude emf is induced between the driver and passenger sides of the truck?

0V

.00178V

.49397V

.07463V

none
Answer #1

1 inch = 0.0254 m

so 61.7 in = 0.0254×61.7 = 1.56718 m

now induced emf by a conductor rod is given by

e = B×L×V

e= 36 ×10^-6 × 1.56718 × 32 = 0.001805 V

nearly equal to option b, 0.00178 V